# Beyond Minimax Optimality: A Subgame Perfect Gradient Method

### v1

@@ 1 Introduction
where we define $\tau^{\text{O}}_{N}$ in (OGM recurrence). [^4] We state the (OGM recurrence) in terms of a sequence $\tau^{\text{O}}_{0},\dots,\tau^{\text{O}}_{N}$. This differs from the original presentation in [14, Equation 6.13] that involves a sequence $\theta_{0},\dots,\theta_{N}$. Equivalence between the two recurrences follows from the change of variables: $\tau^{\text{O}}_{i}=2\theta_{i}^{2}$ for $i=0,\dots,N-1$ and $\tau^{\text{O}}_{N}=\theta_{N}^{2}$. It was later shown in [7] that as long as $d\geq N+2$, there exists a function $f_{\textup{hard}}$ so that for any gradient span first-order method,
@@
While OGM achieves the optimal a priori worst-case guarantee, its guarantees may be overly conservative and its execution may be suboptimal when faced with non-adversarial functions. As a simple example, consider the quadratic function $f(x)=\frac{L}{2}x^{2}$ with initial iterate $x_{0}=1$ (see Figure 1). It is known [15, Theorem 5.1] that OGM’s behavior on this function satisfies $f(x_{N})-f(x_{\star})=\frac{L\left\lVert x_{0}-x_{\star}\right\rVert^{2}}{2\tau^{\text{O}}_{N}}$. However, OGM’s choice of $x_{2},\dots,x_{N}$ are suboptimal, as after seeing just the function and gradient values at $x_{0}$ and $x_{1}$, one can deduce that $x_{\star}=0$ for any $L$-smooth convex function agreeing with this history. We see that OGM provides optimal guarantees on “hard instances”, but fails to improve its guarantees on “easy instances”.
In practical settings, the first-order oracle simply reports first-order information from a fixed function and may at times reveal useful information that would otherwise not be present in an adversarial function. Thus, our goal is to construct a method that exploits useful information, in a dynamic way, to achieve optimal convergence on both “hard instances” and “easy instances” in a formal sense.
@@ 1.2 Equilibrium notions in the minimization game
One strategy for Bob would be to report the function and gradient values from the $L$-smooth function $f_{\textup{hard}}$. This strategy is available as long as $d\geq N+2$. By [7, Theorem 3], Bob can guarantee that
@@
This notion is practically important when the function we are trying to optimize is not adversarially chosen. We may model the non-adversarial nature of typical functions as Bob playing this game “suboptimally”, i.e. playing a strategy which is not part of a Nash Equilibrium for the game. Thus, a strategy for Alice which is part of a Subgame Perfect Equilibrium can be regarded as an algorithm which capitalizes on suboptimal play to the largest extent possible, while still accounting for the possibility that Bob may begin to play adversarially in the future.
@@
Technically, SPGM is not a “strategy” for Alice as SPGM does not assign an action to certain game states, namely those that cannot occur as a result of running SPGM on some first-order oracle. However, because we are primarily concerned with the design of algorithms, such subgames are not relevant to us, as they would not result from choices made by the algorithm. It can be shown via backwards induction that there is a Subgame Perfect Equilibrium $(A,B)$ so that $A$ extends SPGM in the sense that $A$ plays the same action as SPGM in all cases where SPGM is defined, but this is beyond the scope of this paper.
An identical result holds for the setting where history from only the $k$ most recent iterations is stored. In this setting, it is natural to augment the first-order history ${\cal H}$ with a minimal amount of additional information $\tau_{i}\in{\mathbb{R}}$ and $z_{i+1}\in{\mathbb{R}}^{d}$ related to intermediate deductions about the unknown function $f$ made in the $i$th iteration. See Sections 2.2 and 3 for the motivation and definition for $\tau_{i}$ and $z_{i+1}$.
@@ 1.3 Related work
Since the early 2010s, there has been an explosion of work in this area in large part due to the Performance Estimation Program (PEP) framework initiated by Drori and Teboulle [9], Taylor et al. [27]. There, it was shown that the problem of proving a convergence rate for a Fixed-Step First-Order Method (FSFOM) can be rephrased as solving a semidefinite program. Formally, a FSFOM is a first-order method where stepsizes and/or momentum terms are constants that are chosen offline independent of the first-order oracle’s output. This opened the door to the computer-assisted design of first-order methods [9, 13]. By now, a significant list of minimax optimal or conjectured minimax optimal first-order methods have been discovered. In the game-theoretic framework of Section 1.1, minimax optimal corresponds to the best possible payoff for Alice that she can guarantee when playing against an optimal adversary. The Optimized Gradient Method for smooth convex optimization was first observed numerically in [9] and proved analytically in [14]. It was proved to be minimax optimal in [7]. The minimax optimal constant stepsize gradient descent scheme for smooth convex or strongly convex optimization was first observed numerically in [9] and proved in [26, 17]. Conjectured minimax optimal long stepsize gradient descent schemes with partial acceleration were first observed numerically in [13] with upper bounds proved analytically in [1, 2, 11, 12, 28, 29]. A minimax optimal first-order method for strongly convex minimization, known as the Information-Theoretic Exact Method, was derived in [8].
In conjunction with these developments, a significant body of literature has now evolved for producing lower bounds [22, 21, 7, 8]. Our analysis in Section 4 relies heavily on the zero-chain constructions of [8].
We emphasize that all results above are a priori minimax optimal or conjectured minimax optimal. In constrast, the guarantees of SPGM are dynamic minimax optimal. Additionally, the SPGM is not a FSFOM as its stepsizes and momentum terms are chosen dynamically at runtime.
@@ 2.2 Optimized Gradient Method (OGM)
The performance of OGM can be understood in terms of an inductive statement that is maintained throughout the algorithm. We begin the induction by setting $\tau^{\text{O}}_{0}=2$ and partially defining $z_{1}=x_{0}-\frac{2}{L}g_{0}$. Then, the following expression is guaranteed to be nonnegative:
@@ 3 Subgame perfect strategies for the Minimization Game
If $z_{n}\neq x_{0}$, then for any $\alpha\in(0,1)$, the expression in the display above is strictly positive. Thus, (4) is essentially strictly feasible, in the sense of [4, Definition 1.4.2]. This guarantees strong duality holds for (4). ∎
@@
$f(x_{n})-f(x_{\star})\leq\frac{L}{2\tau_{n}}\|x_{0}-x_{\star}\|^{2}$
@@ 4.2 Lower bound guarantees
Conditions ensuring that (14) satisfies (21) for $j\in[n,N-1]$ are given by Drori and Taylor [8, Theorem 3]. In particular, it suffices to show
@@ 5.2 Regression problem instances derived from real data
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\sum_{i=1}^{m}\log(1+\exp(b_{i}\times a_{i}^{\intercal}x))+\frac{1}{2}\|x\|^{2}_{2},$ (28)
using the classification datasets “australian”, “diabetes”, “heart”, “ionosphere”, “splice”, and “sonar”. The performance on these twelve instances is presented in Figure 5.
@@ 6 Conclusion and discussion
This paper offers a framework for analyzing first-order methods beyond the classical notion of minimax optimality. In contrast to minimax optimality, which can be thought of as a strategy in a Nash Equilibrium for the minimization “game”, the notion of a Subgame Perfect Equilibrium requires that a first-order method optimally capitalize on any non-adverarial first-order information. We believe this is a fundamental avenue towards deriving formal beyond-worst-case results in the design of optimization algorithms.

### v2

@@ 1 Introduction
where we define $\tau^{\text{O}}_{N}$ in (OGM recurrence). [^4] We state the (OGM recurrence) in terms of a sequence $\tau^{\text{O}}_{0},\dots,\tau^{\text{O}}_{N}$. This differs from the original presentation in [19, Equation 6.13] that involves a sequence $\theta_{0},\dots,\theta_{N}$. Equivalence between the two recurrences follows from the change of variables: $\tau^{\text{O}}_{i}=2\theta_{i}^{2}$ for $i=0,\dots,N-1$ and $\tau^{\text{O}}_{N}=\theta_{N}^{2}$. It was later shown in [9] that as long as $d\geq N+2$, there exists a function $f_{\textup{hard}}$ so that for any gradient span first-order method,
@@
While OGM achieves the optimal a priori worst-case guarantee, its guarantees may be overly conservative, and its execution may be suboptimal when faced with non-adversarial functions. As a simple example, consider the quadratic function $f(x)=\frac{L}{2}x^{2}$ with initial iterate $x_{0}=1$ (see Figure 1). It is known [20, Theorem 5.1] that OGM’s behavior on this function satisfies $f(x_{N})-f(x_{\star})=\frac{L\left\lVert x_{0}-x_{\star}\right\rVert^{2}}{2\tau^{\text{O}}_{N}}$. However, OGM’s choice of $x_{2},\dots,x_{N}$ is suboptimal, as after seeing just the function and gradient values at $x_{0}$ and $x_{1}$, one can deduce that $x_{\star}=0$ for any $L$-smooth convex function agreeing with this history. We see that OGM provides optimal guarantees on “hard instances”, but fails to improve its guarantees on “easy instances”.
In practical settings, the first-order oracle simply reports first-order information from a fixed function and may, at times, reveal useful information that would otherwise not be present in an adversarial function. Thus, our goal is to construct a method that exploits useful information in a dynamic way to achieve optimal convergence on both “hard instances” and “easy instances” in a formal sense.
@@ 1.2 Equilibrium notions in the minimization game
One strategy for Bob would be to report the function and gradient values from the $L$-smooth function $f_{\textup{hard}}$. This strategy is available as long as $d\geq N+2$. By [9, Theorem 3], Bob can guarantee that
@@
This notion is practically important when the function we are trying to optimize is not adversarially chosen. We may model the non-adversarial nature of typical functions as Bob playing this game “suboptimally”, i.e. playing a strategy that is not part of a Nash Equilibrium for the game. Thus, a strategy for Alice which is part of a Subgame Perfect Equilibrium can be regarded as an algorithm which capitalizes on suboptimal play to the largest extent possible, while still accounting for the possibility that Bob may begin to play adversarially in the future.
@@
Technically, SPGM is not a “strategy” for Alice as SPGM does not assign an action to certain game states, namely those that cannot occur as a result of running SPGM on some first-order oracle. However, because we are primarily concerned with the design of algorithms, such subgames are not relevant to us, as they would not result from choices made by the algorithm. It can be shown via backward induction that there is a Subgame Perfect Equilibrium $(A,B)$ so that $A$ extends SPGM in the sense that $A$ plays the same action as SPGM in all cases where SPGM is defined, but this is beyond the scope of this paper.
An identical result holds for the setting where the history from only the $k$ most recent iterations is stored. In this setting, it is natural to augment the first-order history ${\cal H}$ with a minimal amount of additional information $\tau_{i}\in{\mathbb{R}}$ and $z_{i+1}\in{\mathbb{R}}^{d}$ related to intermediate deductions about the unknown function $f$ made in the $i$th iteration. See Sections 2.2 and 3 for the motivation and definition for $\tau_{i}$ and $z_{i+1}$.
@@ 1.3 Related work
Since the early 2010s, there has been an explosion of work seeking optimal methods in large part due to the Performance Estimation Program (PEP) framework initiated by Drori and Teboulle [12], Taylor et al. [37]. There, it was shown that the problem of proving a convergence rate for a Fixed-Step First-Order Method (FSFOM) can be rephrased as solving a semidefinite program. Formally, a FSFOM is a first-order method where stepsizes and/or momentum terms are constants that are chosen offline independent of the first-order oracle’s output. This opened the door to the computer-assisted design of first-order methods [12, 16]. By now, a significant list of minimax optimal or conjectured minimax optimal first-order methods have been discovered. In the game-theoretic framework of Section 1.1, minimax optimal corresponds to the best possible payoff for Alice that she can guarantee when playing against an optimal adversary.
The Optimized Gradient Method for smooth convex optimization was first observed numerically in [12] and proved analytically in [19]. It was proved to be minimax optimal in [9]. Adaptive versions of this method have been designed using linesearching [10, 35] as well as restarting [22] techniques. An optimized first-order method for strongly convex minimization was derived in [39, 8, 33], culminating in the development of the minimax optimal Information-Theoretic Exact Method of [36]. Optimized methods able to be applied to various constrained settings or settings with proximal terms have been developed as well [38, 21, 18, 4, 17]. The minimax optimal constant stepsize gradient descent scheme for smooth convex or strongly convex optimization was first observed numerically in [12] and proved in [34, 24]. Conjectured minimax optimal long stepsize gradient descent schemes with partial acceleration were first observed numerically in [16] with upper bounds proved analytically in [1, 2, 14, 15, 40, 41].
In conjunction with all of the above algorithmic developments, a significant body of literature has now evolved for producing lower bounds [29, 28, 9, 11]. Our analysis in Section 4 relies heavily on the zero-chain constructions of [11].
We emphasize that all results above are a priori minimax optimal or conjectured minimax optimal. In contrast, the guarantees of SPGM are dynamic minimax optimal. Additionally, the SPGM is not a FSFOM as its stepsizes and momentum terms are chosen dynamically at runtime.
@@ 2.2 Optimized Gradient Method (OGM)
The performance of OGM can be understood in terms of an inductive statement that is maintained throughout the algorithm. The elegant analysis below was first given by [35] and presented clearly in [13, Theorem 4.4] and [33]. We begin the induction by setting $\tau^{\text{O}}_{0}=2$ and partially defining $z_{1}=x_{0}-\frac{2}{L}g_{0}$. Then, the following expression is guaranteed to be nonnegative:
@@ 3 Subgame perfect strategies for the Minimization Game
If $z_{n}\neq x_{0}$, then for any $\alpha\in(0,1)$, the expression in the display above is strictly positive. Thus, (4) is essentially strictly feasible, in the sense of [5, Definition 1.4.2]. This guarantees strong duality holds for (4). ∎
@@
$f(x_{n})-f(x_{\star})\leq\frac{L}{2\tau_{n}}\|x_{0}-x_{\star}\|^{2}+\frac{1}{2}\|g_{n}\|^{2}$
@@ 4.2 Lower bound guarantees
Conditions ensuring that (14) satisfies (21) for $j\in[n,N-1]$ are given by Drori and Taylor [11, Theorem 3]. In particular, it suffices to show
@@ 5.2 Regression problem instances derived from real data
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\sum_{i=1}^{m}\log(1+\exp(b_{i}\cdot a_{i}^{T}x))+\frac{1}{2m}\|x\|^{2}_{2},$ (28)
using normalized data $A,b$ from the classification datasets “australian”, “diabetes”, “heart”, “ionosphere”, “splice”, and “sonar”. The performance on these twelve instances is presented in Figure 5.
@@ 6 Conclusion and discussion
This paper offers a framework for analyzing first-order methods beyond the classical notion of minimax optimality. In contrast to minimax optimality, which can be thought of as a strategy in a Nash Equilibrium for the minimization “game”, the notion of a Subgame Perfect Equilibrium requires that a first-order method optimally capitalize on any non-adversarial first-order information. We believe this is a fundamental avenue towards deriving formal beyond-worst-case results in the design of optimization algorithms.